For the system of linear equations in questions 1
- Determine how many solutions exist
- Use either elimination or substitution to find the solutions (if any)
- Graph the two lines, labeling the x-intercepts, y-intercepts, and points of intersection
3x + 7y = 3 and 12x + 28y = 16

Answer by checkley71(8403)

3x+7y=3
12x+28y=16 multiplying equation 1 by -4 we get
-12x-28y=-64 subtracting these equations we get
-----------
0x+0y=-48 or no intercept
using the slope interception form we get y=-3/7x-3/7 & y=-3/7+4/7 thus the two equations have the same slope -3/7 ergo no interception.